Is(5,-2) a solution of the system?
2x+6y=-2
2x+y=6

Answer by funmath(2933)

Is(5,-2) a solution of the system?
E1) 2x+6y=-2
E2) 2x+y=6
:
Let (x,y)=(5,-2) in E1 and see if both sides are equal:
2(5)+6(-2)=-2
10-12=-2
-2=-2 E1 checks
:
Now let (x,y)=(5,-2) in E2 and see if both sides are equal:
2(5)+(-2)=6
10-2=6
8=6 E2 does NOT check
Therefore (5,-2) is NOT a solution of the system.
Happy Calculating!!!